# Towards Self‑Evolving Benchmarks: Synthesizing Agent Trajectories via Test‑Time Exploration under Validate‑by‑Reproduce Paradigm

### v2

@@ Abstract
Recent advances in large language models (LLMs) and agent system designs have empowered agents with unprecedented levels of capability. However, existing agent benchmarks are showing a trend of rapid ceiling-hitting by newly developed agents, making it difficult to meet the demands for evaluating agent abilities. To address this problem, we propose the Trajectory-based Validated-by-Reproducing Agent-benchmark Complexity Evolution (TRACE) framework. This framework takes an original task from an existing benchmark and encourages agents to freely explore and evolve it into a new task with higher difficulty while recording validatable agent trajectories. The framework proceeds in three stages: (1) evolutionary proposal mining, which provides task evolution proposals through preliminary exploration and divergent thinking; (2) problem formation and free exploration, where proposals are conceptualized into feasible problem candidates and the agents then explore them freely while recording their execution trajectories; and (3) multi-level validation, which ensures that the evolved tasks are accompanied by validatable and reproducible trajectories. Experiments on the GAIA benchmark demonstrate that the TRACE framework consistently enhances task complexity while improving the reliability of correctness through validatable execution trajectories. In addition, our framework can successfully adapt to and improve reasoning datasets represented by AIME-2024. This work marks a paradigm shift from static, manually curated benchmarks to dynamic, self-evolving evaluation systems, providing a sustainable and challenging runway for agent development.
footnotetext: Equal contribution (alphabetical by last name)  Corresponding authorfootnotetext: Work done during an internship at Shanghai Artificial Intelligence Laboratory, supervised by Dongrui Liufootnotetext: This is a work in progress due to methodology refinement and further evaluation
@@ 1 Introduction
The paradigm of artificial intelligence is rapidly shifting towards autonomous agents capable of complex reasoning Comanici et al. (2025); Huang & Yang (2025), planning Huang et al. (2024), and tool utilization Qu et al. (2025); Wang et al. (2024a). This progress is starkly evident in the performance on challenging agent benchmarks Mialon et al. (2023); Jimenez et al. (2023), which were once considered formidable. For instance, on the GAIA benchmark which is designed to test real-world assistant capabilities, top-performing agents have achieved scores exceeding 84% (GAIA Benchmark Team, 2025), rapidly closing the gap with the 92% human baseline. This rapid pace of development, with leaderboards being refreshed almost weekly, signals an urgent challenge: the rapid ceiling-hitting of existing evaluation benchmarks, posing a significant threat to continued progress. When benchmarks become saturated, they lose their ability to differentiate between state-of-the-art agents, obscure true limitations, and may misdirect research efforts towards “overfitting” to a static set of problems rather than pursuing generalizable intelligence. However, the cost of manually creating novel, complex, and reliable tasks is a labor-intensive, time-consuming, and expensive process, which highlights an urgent need for an automated and scalable approach to agent benchmark evolution.
@@
However, evolving agent tasks presents unique challenges not found in conventional domains like mathematical reasoning Hendrycks et al. (2021); Guo et al. (2025) or knowledge-based question answering Yang et al. (2018). Agent tasks are defined by two key characteristics: (1) their procedural nature, which emphasizes complex, multi-step interactions with dynamic real-world environments (e.g., websites, APIs), and (2) their immense diversity, spanning from web navigation to software operation. These characteristics render traditional task evolution methods, such as rule-based parameter mutation or scaling largely ineffective. For instance, rule-based changes Wu & Liu (2025); Wang et al. (2024b) (e.g., altering a specific keyword in a search query) are often too brittle. In a dynamic web environment, such a minor change could break the task’s solvability entirely rather than increasing its reasoning complexity. Similarly, merely scaling up a task Liu et al. (2025) (e.g., asking to book three flights instead of one) often increases repetition, rather than the core cognitive or planning challenges. Thus, a new paradigm is required that moves beyond superficial modifications to fundamentally enhance the procedural, logical and semantic complexity of agent tasks.
@@
To bridge this gap, we introduce the the Trajectory-based Validated-by-Reproducing Agent-benchmark Complexity Evolution (TRACE) framework. Fundamentally, TRACE departs from prior methods that rely on rigid, predefined rules. Instead, it pioneers a novel paradigm, leveraging the generative and agentic capabilities of LLMs as evolutionary engines. This entire process is powered by three specialized agents working in concert: the Evolutionary Proposer, which generates diverse evolution directions through preliminary exploration; the Exploration Executor, which transforms these proposals into feasible problem candidates and conducts free-form exploration while recording execution trajectories; and the Trajectory Validator, which verifies logical coherence and reproducibility, ensuring that the evolved tasks are accompanied by validatable trajectories rather than relying solely on final-answer correctness.
@@
The synergy between these agents facilitates a sophisticated workflow for task evolution. This process commences with the Evolutionary Proposer, which takes an initial task. For instance, “What is the name of Taylor Swift’s debut album?”, and explores its semantic space to identify potential evolutionary forks. Rather than stopping at the answer, it delves into related artifacts like the music video for the single “Teardrops on My Guitar,” generating a portfolio of distinct proposals. These proposals represent diverse vectors for increasing complexity, such as connecting the video’s actor to their filmography or tracing their career across different domains. Subsequently, the Exploration Executor operationalizes these proposals through our core mechanism of proposal-guided trajectory construction. It traverses the original solution path and, upon reaching a relevant state, can dynamically and iteratively integrate compatible proposals. For example, it might first inject a proposal to identify the actor in the video (Tyler Hilton) and the TV series he starred in (One Tree Hill), then layer on a second proposal that pivots to his music career. Through such exploration, the task evolves from a simple single-hop lookup into a multi-step reasoning process with cross-domain dependencies, culminating in a far more sophisticated final question via reverse construction Trinh et al. (2024); Fang et al. (2025); Sun et al. (2024); Gao et al. (2025), such as: “The male lead in the music video for a song from Taylor Swift’s debut album is a multi-talented individual who also starred in a long-running TV series that premiered in 2003. What is the title of this individual’s own debut studio album?”
Importantly, each step of this exploration is recorded as part of a reproducible execution trajectory, which later serves as the basis for validation. Rather than producing only a more complex question, the framework outputs both the evolved task and its accompanying trajectory, ensuring that the added complexity remains transparent, validateable, and reproducible.
This architecture offers an extremely high degree of flexibility and guarantees the integrity of the evolved tasks. The framework’s flexibility stems from two key aspects: first, the clear separation of the Evolutionary Proposer from the Exploration Executor; second, the capacity to evolve tasks along diverse dimensions, including enhancing tool invocation, escalating logical difficulty, or even adopting combinatorial approaches. Such adaptability is not hard-coded but emergent, depending on the bottleneck of the original task, the specific scenario, and the observation and divergent thinking capabilities of the agents themselves. To ensure integrity, the Exploration Executor’s output is not merely the question, but a complete and reproducible trajectory detailing the solution path, tool invocations, executed programs, and their results. This trajectory is then submitted to the Trajectory Validator for a rigorous verification. The Validator first audits the logical coherence of the problem-solving path. Second, it re-executes each step individually to confirm that the outputs match the Executor’s records, ensuring verifiability and reproducibility. This process also inherently examines the task’s solvability and the uniqueness of its answer, forming a robust quality assurance loop for generating certifiably complex and sound benchmarks.
@@
  • We propose TRACE (Trajectory-based Validated-by-Reproducing Agent-benchmark Complexity Evolution), a self-evolving benchmark framework that encourages agent exploration and records execution trajectories as first-class artifacts, ensuring transparency and reproducibility in task evolution.
  • We empirically validate Trace on challenging benchmarks such as GAIA, demonstrating that it consistently produces tasks of higher difficulty, on which prominent agent systems exhibit significant performance degradation shown in Figure 1, thereby validating the effectiveness of our approach.
  • Beyond “one-more-hop” edits, our experiments reveal a From Seed to Spark pattern: under our TRACE framework, the model autonomously explores and evolves problems that can shift into entirely different capability domains (e.g., from retrieval to math+coding), thereby substantially increasing the diversity of evolved tasks. The evolved tasks exhibit greater task diversity and require deeper reasoning depth, providing a robust methodology for rigorously evaluating and advancing future AI agents.
@@ 2 Related Work
In recent years, a variety of static agent benchmarks have been developed to rigorously evaluate the autonomous problem-solving capability of LLM-based agents. GAIA (Mialon et al., 2023) collects 466 human-centric questions requiring reasoning, multimodal understanding, web browsing, and tool use proficiency. The USACO benchmark (Shi et al., 2024) adapts 307 problems from the USA Computing Olympiad(USACO) with high-quality unit tests, reference solutions, and official analysis to assess language models on complex algorithmic reasoning and competitive programming challenges. MLE-bench (Chan et al., 2024) curates 75 real-world ML engineering competitions from Kaggle, spanning data preparation, model training, and experiment management, measuring agents’ end-to-end machine learning capabilities against human baselines. SWE-bench (Jimenez et al., 2023) compiles 2,294 genuine software engineering issues from GitHub repositories, tasking agents with generating code patches to resolve real bugs under authentic development contexts.
Beyond answer-verified benchmarks, a complementary line of work evaluates how agents solve tasks, not merely whether they succeed. WebArena (Zhou et al., 2023) provides self-hosted websites with executable scoring and reproducible trajectory replay, enabling both step- and task-level analyses. Mind2Web (Deng et al., 2023) contributes human-demonstrated trajectories on real websites with fine-grained metrics such as element accuracy and step success rate. Its multimodal extension (Pahuja et al., 2025) further annotates trajectories with multimodal signals and reports step-wise metrics, highlighting process sensitivity in open-ended tasks. Existing works are essentially static and pre-defined. these benchmarks face rapid saturation under the exponential progress of agents: tasks once considered formidable are increasingly solved with near-human accuracy, diminishing their power to differentiate emerging models. Moreover, the collection and curation of such datasets incur substantial human labor cost, making it impractical to refresh benchmarks at a frequency that matches model iteration cycles. As a result, existing benchmarks risk being quickly “outpaced” by the systems they aim to evaluate, underscoring the need for a scalable, self-evolving alternative.
@@
Recent advances in benchmark evolution aim to address the rapid saturation of static evaluation datasets. Benchmark Self-Evolving (Wang et al., 2024b) and AutoEvoEval (Wu & Liu, 2025) represent predefined atomic operation-based approaches in NLP reasoning and multi-choice QA. The former applies six reframing operations to modify contexts or questions, while the latter employs 22 interpretable operations for close-ended tasks, enabling multi-round compositions that mainly test robustness via structural and semantic perturbations. Both operate through surface-level transformations, focusing on robustness rather than evolving underlying task complexity.
AdamMeme (Chen et al., 2025) adopts an agent-based framework to iteratively update meme datasets, probing multimodal LLMs’ reasoning about harmfulness. Through multi-agent collaboration, it refines challenging meme samples, exposing weaknesses in mLLM interpretations. However, its methodology is confined to this single subdomain and does not generalize. EvoCodeBench (Li et al., 2024) builds an automated pipeline that ingests real-world code repositories and updates the benchmark at fixed intervals (e.g., every six months) to prevent data leakage. Unlike methods leveraging model feedback, it relies solely on curated data and rigid schedules.
TRACE pursues benchmark evolution through a fundamentally different paradigm. Rather than relying on predefined operations or rigid update schedules, it leverages test-time exploration as a generative mechanism. Agents autonomously discover and construct harder tasks, producing both the evolved problem and its execution trajectory. Trajectories undergo strict validation, ensuring tasks are more complex, reproducible, and answer-validatable. TRACE enables self-evolving benchmark without human intervention, introducing a trajectory-aware paradigm for scalable evaluation.
@@ 3.1 Agentic Workflow as a DAG
In this section, we represent an agentic workflow $W$ as a directed acyclic graph (DAG) $G=(\mathcal{S},\mathcal{E})$, where the node set $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{N}\}$ corresponds to an ordered sequence of discrete LLM-invoking steps, and the edge set $\mathcal{E}\subseteq\mathcal{S}\times\mathcal{S}$ encodes data dependencies and control-flow constraints—specifically, an edge $(S_{i},S_{j})\in\mathcal{E}$ signifies that the output of step $S_{i}$ is required as input for step $S_{j}$. Each node can be abstracted as the quadruple below:
@@
After execution, the context is updated as $c_{i}=c_{i-1}\oplus(a_{i},o_{i})$, and edges $(S_{i},S_{j})$ arise when artifacts from $S_{i}$—typically $o_{i}$ or the updated context $c_{i}$—are consumed by $S_{j}$; tool-free steps are a special case where $a_{i}$ is a user-facing message and $o_{i}$ may be empty.
@@ 3.3 Problem Statement
Here, $\mathcal{T}_{\text{validatable}}$ denotes the set of trajectories satisfying the validity conditions defined above. The TRACE framework addresses this objective through a three-stage pipeline: proposal mining, proposal-guided exploration, and multi-level validation.
@@ 4.1 Stage 1: Evolutionary Proposal Mining
The Evolution Proposer takes the description of an original task from an existing dataset, optionally along with potential solution paths and answers as input, then outputs multiple evolutionary proposals with diversity. This is operationalized by an LLM agent acting as an expert intelligent agent task designer. The agent is prompted with a comprehensive set of evolutionary strategies, encouraging it to increase difficulty by lengthening the required evidence chain, complicating tool use, targeting specialized domains, or escalating the core reasoning demands. To steer this creative process toward generating sound and innovative challenges, the agent’s behavior is governed by a set of core Guiding Principles. These principles grant the agent the autonomy to think divergently and even pivot to entirely new scenarios if the original task has limited potential for evolution. Crucially, they also enforce that all proposed modifications must lead to deterministic and validatable solutions, preventing ambiguity. Ultimately, the Proposer consolidates its reasoning into a set of diverse and actionable evolutionary proposals, each providing a clear, imperative instruction for how the task should be modified.

### v3

# Towards Self-Evolving Agent Benchmarks: Validatable Agent Trajectory via Test-Time Exploration

@@ Abstract
Recent advances in large language models (LLMs) and agent system designs have empowered agents with unprecedented levels of capability. However, existing agent benchmarks are showing a trend of rapid ceiling-hitting by newly developed agents, making it increasingly difficult to meet the demands of evaluating agent abilities. To address this problem, we propose the Trajectory-based Validated-by-Reproducing Agent-benchmark Complexity Evolution (TRACE) framework. This framework takes an original task from an existing benchmark and encourages agents to freely explore and evolve it into a new task with higher difficulty while recording the corresponding execution trajectories. The framework proceeds in three stages: (1) evolutionary proposal mining, which generates task evolution proposals through preliminary exploration and divergent thinking; (2) problem construction via free exploration, where proposals are instantiated into concrete problem instances through agent exploration, with execution trajectories recorded along the process; and (3) multi-level validation, which ensures that the evolved tasks are accompanied by reproducible and logically coherent trajectories. Experiments on the GAIA benchmark demonstrate that the TRACE framework consistently enhances task complexity while improving correctness reliability through trajectory-level validation. In addition, our framework can successfully adapt to and improve reasoning benchmarks such as AIME-2024. This work marks a paradigm shift from static, manually curated benchmarks to dynamic, self-evolving evaluation systems, providing a sustainable and challenging foundation for agent development.
footnotetext: Equal contribution. Work done during an internship at Shanghai Artificial Intelligence Laboratory, supervised by Dongrui Liufootnotetext: Corresponding author
@@ 1 Introduction
The paradigm of artificial intelligence is rapidly shifting towards autonomous agents capable of complex reasoning (Comanici et al., 2025; Huang and Yang, 2025), planning (Huang et al., 2024), and tool utilization (Qu et al., 2025; Wang et al., 2024a). This progress is starkly evident in the performance on challenging agent benchmarks (Mialon et al., 2023; Jimenez et al., 2023), which were once considered formidable. For instance, on the GAIA benchmark which is designed to test real-world assistant capabilities, top-performing agents have achieved scores exceeding 90% (GAIA Benchmark Team, 2025), rapidly closing the gap with the human baseline. This rapid pace signals an urgent challenge: existing benchmarks are becoming saturated, diminishing their ability to differentiate state-of-the-art agents and risking progress being driven by overfitting to static test sets rather than generalizable intelligence. However, the cost of manually creating novel, complex, and reliable tasks is a labor-intensive, time-consuming, and expensive process, which highlights an urgent need for an automated and scalable approach to agent benchmark evolution.
@@
However, evolving agent tasks presents unique challenges not found in conventional domains like mathematical reasoning (Hendrycks et al., 2021; Guo et al., 2025) or knowledge-based question answering (Yang et al., 2018). Agent tasks are defined by two key characteristics: (1) their procedural nature, which emphasizes complex, multi-step interactions with dynamic, scalable real-world environments (e.g., websites, APIs) (Huang et al., 2025; Liu et al., 2025a; Gao et al., 2025a), and (2) their immense diversity, spanning from web navigation to software operation. These characteristics render traditional task evolution methods, such as rule-based parameter mutation or scaling largely ineffective. For instance, rule-based changes (Wu and Liu, 2025; Wang et al., 2024b) (e.g., altering a specific keyword in a search query) are often insufficiently robust. In a dynamic web environment, such a minor change could break the task’s solvability entirely rather than increasing its reasoning complexity. Similarly, merely scaling up a task (Liu et al., 2025b) (e.g., asking to book three flights instead of one) often increases repetition, rather than the core cognitive or planning challenges. Thus, a new paradigm is required that moves beyond superficial modifications to fundamentally enhance the procedural, logical and semantic complexity of agent tasks.
@@
To bridge this gap, we introduce TRACE (Trajectory-based Validated-by-Reproducing Agent-benchmark Complexity Evolution). Departing from rigid, rule-based heuristics, TRACE leverages the agentic capabilities of LLMs to drive task evolution through three specialized components. The Evolutionary Proposer identifies diverse evolution directions via preliminary exploration; the Exploration Executor navigates real-world environments to materialize these proposals into recorded execution trajectories; and the Trajectory Validator enforces logical coherence and reproducibility. This ensures that evolved tasks are grounded in verifiable solution paths rather than relying solely on final-answer correctness.
@@
The coordination between these agents enables a structured workflow for task evolution. This process begins with the Evolutionary Proposer, which analyzes task bottlenecks and explores its semantic space to identify evolutionary forks. Consider the question: “What is the name of Taylor Swift’s debut album?” Rather than stopping at the direct answer, it examines related artifacts, such as the music video for the single Teardrops on My Guitar, and generates multiple evolution proposals. These proposals introduce additional reasoning dependencies, such as linking the video’s actor to their filmography or tracing their cross-domain career trajectory. Subsequently, the Exploration Executor materializes these proposals through proposal-guided trajectory construction. It traverses the original solution path and integrates compatible proposals at appropriate intermediate states. For example, it may first inject a proposal to identify the actor in the video (Tyler Hilton) and the TV series he starred in (One Tree Hill), then introduce a second proposal that pivots to his music career. This process transforms a single-hop lookup into a multi-step reasoning task with cross-domain dependencies, culminating in a more complex final question (Trinh et al., 2024; Fang et al., 2025; Sun et al., 2024; Gao et al., 2025b), such as: “The male lead in the music video for a song from Taylor Swift’s debut album is a multi-talented individual who also starred in a long-running TV series that premiered in 2003. What is the title of this individual’s own debut studio album?” More broadly, this bottleneck-aware proposal and injection process extends beyond information-seeking tasks to reasoning-intensive, mathematical, and coding problems, demonstrating the generality of the workflow. Importantly, each exploration step is recorded as an execution trajectory that serves as the basis for validation, ensuring that the introduced complexity remains transparent and verifiable.
This architecture balances flexibility and rigor in task evolution. Flexibility arises from the separation of the Evolutionary Proposer and the Exploration Executor, enabling evolution along multiple dimensions, such as deeper tool integration, increased logical dependency, or combinatorial composition, without reliance on hard-coded templates. To maintain rigor, the Executor outputs the evolved task together with its execution trajectory, including reasoning steps, tool invocations, and observed results. The Trajectory Validator then audits logical coherence and re-executes each step to confirm reproducibility. This validation loop ensures that generated benchmarks are complex, solvable, and grounded in executable evidence.
In essence, TRACE models the workflow of a human benchmark designer while replacing manual curation with an autonomous evolution process. Unlike static rule-based approaches, it enables agents to explore real-world environments (e.g., the live internet) as structured task substrates. This exploration is guided by bottleneck analysis: by diagnosing limitations of a seed task such as shallow tool usage or limited reasoning depth, the framework steers evolution toward targeted structural augmentation. As a result, TRACE evolves tasks into tool-integrated, multi-step challenges that better reflect real-world problem-solving dynamics, repositioning the agent from a passive solver to an active benchmark constructor.
@@
  • We propose TRACE (Trajectory-based Validated-by-Reproducing Agent-benchmark Complexity Evolution), a benchmark self-evolving framework that encourages agent exploration and records execution trajectories as first-class artifacts, ensuring transparency and reproducibility in task evolution.
  • We empirically validate TRACE on general-purpose agent benchmarks such as GAIA as well as reasoning benchmarks such as AIME-2024, demonstrating that it consistently produces tasks of higher difficulty, on which prominent agent systems exhibit significant performance degradation shown in Figure 1, thereby validating the effectiveness of our approach.
  • Beyond “one-more-hop” edits, our experiments reveal a From Seed to Spark pattern: under our TRACE framework, the model autonomously explores and evolves problems that can shift into entirely different capability domains (e.g., from retrieval to math and coding), thereby substantially increasing the diversity of evolved tasks. The evolved tasks exhibit greater task diversity and require deeper reasoning depth, providing a robust methodology for rigorously evaluating and advancing future agents.
@@ 2 Related Work
A range of agent benchmarks have been proposed to evaluate LLM-based agents under realistic problem-solving settings. GAIA (Mialon et al., 2023) focuses on human-centric questions that require reasoning, multimodal understanding, web browsing, and tool use. USACO (Shi et al., 2024) adapts competitive programming problems with unit tests and reference solutions to assess algorithmic reasoning. MLE-bench (Chan et al., 2024) evaluates end-to-end ML engineering by grounding tasks in Kaggle-style competitions spanning data preparation, model training, and experiment management. SWE-bench (Jimenez et al., 2023) measures software engineering ability through real GitHub issues that require generating patches under authentic repository contexts.
Beyond answer-verified benchmarks, several works emphasize process-level evaluation by analyzing agent trajectories. WebArena (Zhou et al., 2023) provides self-hosted websites with executable scoring and trajectory replay for reproducible step- and task-level analyses. Mind2Web (Deng et al., 2023) contributes human demonstrations on real websites with fine-grained metrics (e.g., element accuracy and step success), and its multimodal extension (Pahuja et al., 2025) augments trajectories with multimodal signals and step-wise evaluation.
Despite their breadth, these benchmarks are largely static: tasks and environments are pre-defined and updated infrequently. As agent capability improves, static benchmarks can saturate and become costly to refresh at a cadence aligned with model iteration, motivating scalable approaches to benchmark evolution.
@@
Benchmark evolution has been explored to mitigate saturation of static datasets. Benchmark Self-Evolving (Wang et al., 2024b) and AutoEvoEval (Wu and Liu, 2025) apply predefined atomic operations to reformulate reasoning and close-ended QA tasks via structural or semantic perturbations, primarily targeting robustness rather than increasing procedural task complexity. AdamMeme (Chen et al., 2025b) adopts multi-agent refinement to iteratively update meme datasets for probing harmfulness-related reasoning, but is specialized to this domain. EvoCodeBench (Li et al., 2024) updates code benchmarks by periodically ingesting new repositories to reduce leakage, relying on curated data streams and fixed schedules rather than model-driven exploration.
In contrast, TRACE evolves agent tasks via test-time exploration in real-world environments. Instead of predefined edits or scheduled refreshes, agents autonomously construct harder tasks and output their execution trajectories, which are subsequently validated for logical coherence and reproducibility. This trajectory-aware evolution supports scalable benchmark updating while preserving solvability and verifiability.
@@ 3.1 Agentic Workflow as a DAG
In this section, we represent an agentic workflow $W$ as a directed acyclic graph (DAG) $G=(\mathcal{S},\mathcal{E})$, where the node set $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{N}\}$ corresponds to an ordered sequence of discrete LLM-invoking steps, and the edge set $\mathcal{E}\subseteq\mathcal{S}\times\mathcal{S}$ encodes data dependencies and control-flow constraints. Specifically, an edge $(S_{i},S_{j})\in\mathcal{E}$ signifies that the output of step $S_{i}$ is required as input for step $S_{j}$. Each node can be abstracted as the quadruple below:
@@
After execution, the context is updated as $c_{i}=c_{i-1}\oplus(a_{i},o_{i})$, and edges $(S_{i},S_{j})$ arise when artifacts from $S_{i}$, typically $o_{i}$ or the updated context $c_{i}$, are consumed by $S_{j}$; tool-free steps are a special case where $a_{i}$ is a user-facing message and $o_{i}$ may be empty.
@@ 3.3 Problem Statement
Here, $\mathcal{T}_{\text{validatable}}$ denotes the set of trajectories satisfying the validity conditions defined above. In practice, we operationalize $\mathrm{Difficulty}(\cdot)$ via a trajectory-informed bottleneck assessment inspired by theory-of-mind (Chen et al., 2025a): given the solution trajectories of the seed and evolved tasks, the difficulty judge estimates which task would impose a larger intrinsic bottleneck for a solver with comparable planning, reasoning, and tool-use capabilities (including access to a similar tool set), potentially requiring additional attempts or a non-trivial conceptual insight to resolve. The TRACE framework addresses this objective through a three-stage pipeline: proposal mining, proposal-guided exploration, and multi-level validation.
@@ 4.1 Stage 1: Evolutionary Proposal Mining
The Evolution Proposer takes the description of an original task from an existing dataset, optionally along with potential solution paths and answers, as input. Rather than directly generating modifications, it first performs a bottleneck-aware pre-exploration stage. In this stage, the agent analyzes the original task and its solution trajectory to identify which agent capabilities the task primarily evaluates (e.g., planning, reasoning, or tool use) and where a solver is most likely to encounter intrinsic difficulty. Based on this bottleneck analysis, the agent then probes the task’s semantic space to uncover dimensions that can be further strengthened such as extending evidence chains, increasing tool interaction complexity, or deepening reasoning requirements, thereby formulating targeted evolution proposals.
Based on this preliminary diagnosis, the proposer then synthesizes multiple evolutionary proposals. It is prompted with a comprehensive set of evolutionary strategies that serve as directional guidance rather than rigid transformation rules. These strategies encourage the agent to lengthen evidence chains, complicate tool interactions, target specialized domains, or escalate reasoning demands, while preserving high degrees of creative freedom.
To ensure that exploration remains principled rather than arbitrary, the proposer is governed by a set of core guiding principles. These principles grant the autonomy to think divergently and even pivoting to new but semantically grounded scenarios when appropriate, while enforcing that all proposed modifications must yield deterministic and validatable solutions. Ultimately, the proposer consolidates its reasoning into a diverse set of actionable evolutionary proposals, each formulated as a clear and imperative instruction for how the seed task should be transformed.